# Efficient Algorithms for Low Tubal Rank Tensor Approximation with Applications to Image Compression, Super-Resolution and Deep Learning

### v1

@@ 9      $[\underline{\bf Q}_{c},\sim]={\rm T\operatorname{-}QR}(\underline{\bf Y}_{c})$;
12 $\underline{\bf Z}=\widehat{\underline{\bf R}}^{-1}*\widehat{\underline{\bf Q}}*{\underline{\bf Y}}^{T}_{r}$;
@@
$\underline{\bf U}=\underline{\bf Q}*\widehat{\underline{\bf U}}$;
@@ 5 Proposed randomized fixed-precision algorithms
Let $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $\underline{\bf\Omega}\in\mathbb{R}^{I_{2}\times k\times I_{3}}$ be a random tensor $\underline{\bf Y}=\underline{\bf X}*\underline{\bf\Omega},\,\underline{\bf W}=\underline{\bf X}^{T}*\underline{\bf Y}$ and the economic T-SVD of $\underline{\bf Y}$ is $\underline{\bf Y}=\widehat{\underline{\bf U}}*\widehat{\underline{\bf S}}*\widehat{\underline{\bf V}}^{T}$. Setting
@@
gives an approximation $\underline{\bf Q}*\underline{\bf B}$ to $\underline{\bf X}$ that has the same accuracy of … and
@@
where $\underline{\bf H}_{1}=\underline{\bf H}(:,:,1)$ is the first frontal slie of $\underline{\bf H}=\underline{\bf W}*\underline{\bf W}^{T}*(\underline{\bf Y}*\underline{\bf Y})^{-1}.$
@@
Now, using the fact that $\|\underline{\bf B}\|_{F}^{2}={\rm trace}(\underline{\bf G}_{1})=\underline{\bf G}(:,:,1)$ where $\underline{\bf G}_{1}$ is the first frontal slice of $\underline{\bf G}=\underline{\bf B}^{T}*\underline{\bf B}$, we have
@@ 6 Experimental results
$\underline{\bf X}_{clean}={\rm randn}(I_{1},R,I_{3})*{\rm randn}(R,I_{1},I_{3}),$ (19)
and add a noise term to it to generate a noisy tensor $\underline{\bf X}_{clean}+\delta\frac{\underline{\bf Y}}{\|\underline{\bf Y}\|_{F}}\|\underline{\bf X}_{\epsilon}\|_{F}$ where $\underline{\bf Y}$ is a standard Gaussian tensor of the same size as the original tensor $\underline{\bf X}$. We set $r=50,\,\delta=10^{-3}$ and assume $I_{1}=I_{2}=I_{3}$ in all our simulations. So, the tubal rank of the tensor $\underline{\bf X}_{clean}$ is approximately equal to $50$, and we checked how accurate the proposed fixed-precision algorithms are to estimate this tubal rank. For a given error tolerance $\epsilon=10^{-5}$ and block size $b=100$, the proposed fixed-precision algorithms 10-11 and the baselines methods including algorithm 9 and the truncated T-SVD are applied to the mentioned data tensor $\underline{\bf X}_{clean}$. Note that Algorithms 9-11 can estimate the tubal rank automatically, while the truncated T-SVD requires the tubal rank as an input. Because of this issue, as mentioned, we first estimated the tubal rank using the fixed-precision algorithms 9-11, and then the estimated tubal rank was used in the truncated T-SVD. Our first observation was that all fixed-precision algorithms estimated the tubal rank correctly, but the CPU time of the proposed algorithms was relatively lower than the baseline Algorithm 9.
Also, for the tubal rank $R=50$, we applied the truncated T-SVD for the tensor $\underline{\bf X}_{clean}$. Here, the running time of the truncated T-SVD was significant compared to the proposed fixed-precision algorithms, while its accuracy was almost the same as that of the proposed randomized algorithms. We also observe that as the tensor size increases, the proposed algorithms’ running times are much less than the baselines. These observations convinced us that the proposed fixed-precision algorithms are faster and more efficient than the baselines.
@@
(Image Compression) We examine the effectiveness of the suggested randomized single-pass algorithms for the image compression problem in this experiment. We used the Kodak dataset  [^3] https://r0k.us/graphics/kodak/ and considered four images, ”Kodim15”, ”Kodim17”, ”Kodim18,” and ”Kodim23” in our simulations. The first two images are of size $512\times 768\times 3$ while the second two images are of size $768\times 512\times 3$. We applied our proposed algorithms and compared them with the baselines single-pass T-CUR [37] and tensor Sketch [32]. In our simulations we set sketches $L=350,\,K=350,\,H=100$ and $R=30$. The reconstructed images with corresponding PSNR and relative errors are reported in Figure 3. The results demonstrate that our proposed randomized single-pass algorithms provided better results than the single-pass T-CUR and the tensor Sketch algorithms. It is interesting to report that we found experimentally that when we set $K=L$, all algorithms except the proposed randomized single-pass Algorithms 6-8 were unstable and provided poor results. In this sense, our algorithm 7 can be more reliable. Better results are achieved as soon as the sketch parameters are different $L\neq K$. The next numerical finding, which we should point out, is that for the case of $L<K$, the results were poor, and in all experiments, this was the case. This instability is due to the ill-conditioning of the underlying least-squares problem, which needs to be solved. More precisely, the Moore-Penrose of the coefficient tensor, according to the least-squares problem, is poorly computed, and the final results are not satisfactory. However, the proposed randomized single-pass Algorithms 6-8 are robust because truncated T-SVD in lines (6-7) act as regularizers and improve the numerical efficiency. This illustration amply demonstrates how the suggested approach is more reliable and capable of producing superior results.
@@
(Video compression) In this example, we study the performance of the proposed randomized single-pass algorithms for the video compression task. We mainly use the videos datasets “Foreman” and “News” [^4] http://trace.eas.asu.edu/yuv/. The size of these videos are third-order tensors of size $144\times 176\times 300$. First, we tested the efficiency of the single-pass algorithms for low tubal rank approximations for the mentioned videos using the sketch parameters $K=90,L=90,H=20$, and $R=20$. The PSNR of all frames of the Foreman and News videos achieved by the proposed single-pass algorithms and the baselines are reported in Figures 4 and 5, respectively. Also, the reconstructed images of some frames of the mentioned videos are displayed in figures 6 and 7, respectively. We see that the proposed Algorithms 6-8 are robust in terms of the sketch sizes used, and this example illustrates the reliability of the proposed randomized single-pass algorithms for compressing videos.
@@
We considered the YOLO3 and applied this network to detect objects in degraded images. These results are shown in Figure 9 (first row). As can be seen, the network for the dog image only detected one object (cat), which is also not correct. The network also detects two objects for the horse image: one horse and one giraffe. It failed to detect two additional horses and also incorrectly detected a horse as a giraffe. In contrast, as a preprocessing stage, our completion technique first quickly processes the degraded image, and then the YOLO3 is applied to it to detect the objects inside the image. After the fast preprocessing using our techniques, the network detected three objects for the dog image: a bicycle, a dog, and a truck with very precise bounding boxes. Also, for the horse’s image, the network detected four horses. The results clearly show the efficiency and feasibility of the proposed completion procedure for stabilizing YOLO3 against pixel removal.

### v2

@@ 3 Apply the T-QR applied to $\underline{\bf Y}$ and obtain the tensor $\underline{\bf Q}$;
1 $\underline{\bf\Omega}_{1}={\rm randn}(I_{2},K+R,I_{3}),\,\,\,\,$ $\underline{\bf\Omega}_{2}={\rm randn}(I_{1},L+R,I_{3})$;
@@ 9      $[\underline{\bf Q}_{c},\sim]={\rm T\operatorname{-}QR}(\underline{\bf Y}_{c})$;
12 $\underline{\bf Z}=\widehat{\underline{\bf R}}^{-1}*\widehat{\underline{\bf Q}}^{T}*{\underline{\bf Y}}^{T}_{r}$;
@@
$\underline{\bf U}=\widehat{\underline{\bf Q}}*\widehat{\underline{\bf U}}$;
@@ 5 Proposed randomized fixed-precision algorithms
Let $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $\underline{\bf\Omega}\in\mathbb{R}^{I_{2}\times k\times I_{3}}$ be a random tensor, $\underline{\bf Y}=\underline{\bf X}*\underline{\bf\Omega},\,\underline{\bf W}=\underline{\bf X}^{T}*\underline{\bf Y}$, and the economic T-SVD of $\underline{\bf Y}$ is $\underline{\bf Y}=\widehat{\underline{\bf U}}*\widehat{\underline{\bf S}}*\widehat{\underline{\bf V}}^{T}$. Setting
@@
gives an approximation $\underline{\bf Q}*\underline{\bf B}$ to $\underline{\bf X}$ that has the same accuracy of the basic randomized algorithms and
@@
where $\underline{\bf H}_{1}=\underline{\bf H}(:,:,1)$ is the first frontal slice of $\underline{\bf H}=\underline{\bf W}*\underline{\bf W}^{T}*(\underline{\bf Y}*\underline{\bf Y})^{-1}.$
@@
Note since $\underline{\bf Q}$ is the orthonormalization of $\underline{\bf X}*\underline{\bf\Omega}$, this approximation $\underline{\bf Q}*\underline{\bf B}$ performs with the same accuracy as the basic randomized algorithms without power iteration and oversampling [25]. Now, using the fact that $\|\underline{\bf B}\|_{F}^{2}={\rm trace}(\underline{\bf G}_{1})=\underline{\bf G}(:,:,1)$ where $\underline{\bf G}_{1}$ is the first frontal slice of $\underline{\bf G}=\underline{\bf B}^{T}*\underline{\bf B}$, we have
@@ 6 Experimental results
$\underline{\bf X}_{\rm clean}={\rm randn}(I_{1},R,I_{3})*{\rm randn}(R,I_{1},I_{3}),$ (19)
and add a noise term to it to generate a noisy tensor $\underline{\bf X}_{\rm perturb}=\underline{\bf X}_{\rm clean}+\delta\frac{\underline{\bf Y}}{\|\underline{\bf Y}\|_{F}}\|\underline{\bf X}_{\epsilon}\|_{F}$ where $\underline{\bf Y}$ is a standard Gaussian tensor of the same size as the original tensor $\underline{\bf X}$. We set $r=50,\,\delta=10^{-3}$ and assume $I_{1}=I_{2}=I_{3}$ in all our simulations. So, the tubal rank of the tensor $\underline{\bf X}_{\rm perturb}$ is approximately equal to $50$, and we checked how accurate the proposed fixed-precision algorithms are to estimate this tubal rank. For a given error tolerance $\epsilon=10^{-5}$ and block size $b=100$, the proposed fixed-precision algobaseline methods,the baselines methods including algorithm 9 and the truncated T-SVD are applied to the mentioned data tensor $\underline{\bf X}_{\rm perturb}$. Note that Algorithms 9-11 can estimate the tubal rank automatically, while the truncated T-SVD requires the tubal rank as an input. Because of this issue, as mentioned, we first estimated the tubal rank using the fixed-precision algorithms 9-11, and then the estimated tubal rank was used in the truncated T-SVD. Our first observation was that all fixed-precision algorithms estimated the tubal rank correctly, but the CPU time of the proposed algorithms was relatively lower than the baseline Algorithm 9.
Also, for the tubal rank $R=50$, we applied the truncated T-SVD to the tensor $\underline{\bf X}_{\rm perturb}$. Here, the running time of the truncated T-SVD was significant compared to the proposed fixed-precision algorithms, while its accuracy was almost the same as that of the proposed randomized algorithms. We also observe that as the tensor size increases, the proposed algorithms’ running times are much less than the baselines. These observations convinced us that they are faster and more efficient than the baselines.
@@
(Image Compression) We examine the effectiveness of the suggested randomized single-pass algorithms for the image compression problem in this experiment. We used the Kodak dataset  [^3] https://r0k.us/graphics/kodak/ and considered four images, ”Kodim15”, ”Kodim17”, ”Kodim18,” and ”Kodim23” in our simulations. The first two images are of size, $512\times 768\times 3$ while the second two images are of size $768\times 512\times 3$. We applied our proposed algorithms and compared them with the baselines single-pass T-CUR [37] and tensor sketch [32]. In our simulations we set sketches $L=350,\,K=350,\,H=100$ and $R=30$. The reconstructed images with corresponding PSNR and relative errors are reported in Figure 3. The results demonstrate that our proposed randomized single-pass algorithms provided better results than the single-pass T-CUR and the tensor sketch algorithms. It is interesting to report that we found experimentally that when we set $K=L$, all algorithms except the proposed randomized single-pass Algorithms 6-8 were unstable and provided poor results. In this sense, our algorithm 7 can be more reliable. Better results are achieved as soon as the sketch parameters are different $L\neq K$. The next numerical finding, which we should point out, is that for the case of $L<K$, the results were poor, and in all experiments, this was the case. This instability is due to the ill-conditioning of the underlying least-squares problem, which needs to be solved. More precisely, the Moore-Penrose of the coefficient tensor, according to the least-squares problem, is poorly computed, and the final results are not satisfactory. However, the proposed randomized single-pass Algorithms 6-8 are robust because truncated T-SVD in lines acts as a regularizer and improves the numerical efficiency. This illustration amply demonstrates how the suggested approach is more reliable and capable of producing superior results.
@@
(Video compression) In this example, we study the performance of the proposed randomized single-pass algorithms for the video compression task. We mainly use the video datasets “Foreman” and “News” [^4] http://trace.eas.asu.edu/yuv/. The size of these videos is third-order tensors of size $144\times 176\times 300$. First, we tested the efficiency of the single-pass algorithms for low tubal rank approximations for the mentioned videos using the sketch parameters $K=90,L=90,H=20$, and $R=20$. The PSNR of all frames of the Foreman and News videos achieved by the proposed single-pass algorithms and the baselines are reported in Figures 4 and 5, respectively. Also, the reconstructed images of some frames of the mentioned videos are displayed in figures 6 and 7, respectively. We see that the proposed Algorithms 6-8 are robust in terms of the sketch sizes used, and this example illustrates the reliability of the proposed randomized single-pass algorithms for compressing videos.
@@
We considered the YOLO3 and applied this network to detect objects in degraded images. These results are shown in Figure 9 (first row). As can be seen, the network for the dog image only detected one object (cat), which is also not correct. The network also detects two objects for the horse image: one horse and one giraffe. It failed to detect two additional horses and also incorrectly detected a horse as a giraffe. In contrast, as a preprocessing stage, our completion technique first quickly processes the degraded image, then the YOLO3 is applied to it to detect the objects inside the image. After the fast preprocessing using our techniques, the network detected three objects for the dog image: a bicycle, a dog, and a truck with very precise bounding boxes. Also, for the horse’s image, the network detected four horses. The results clearly show the efficiency and feasibility of the proposed completion procedure for stabilizing YOLO3 against pixel removal.